### v1

# Spatially Resolved Kinematics of Post-starburst Galaxies

Abstract
Star-forming galaxies can reach quiescence via rapid transition through merger-triggered starbursts that consequently affect both their kinematics and star formation rates. In this work, we analyze the spatially-resolved kinematics of 92 post-starburst galaxies (PSBs) with data from the MaNGA survey and place them in context with early-type galaxies (ETGs) to study the impact of merger history on galaxy kinematics. We measure the specific angular momentum to characterize them as slow or fast rotators. We find that the MaNGA PSB sample has $5\%$ slow rotators, which is less than the $14\%$ slow rotators in the ATLAS3D ETG sample and $22\%$ slow rotators in the MaNGA ETGs. The PSBs have lower specific angular momentum than the star-forming galaxies (SFs) and a greater fraction of slow rotators. This implies that for the PSBs to evolve into the ETGs, they must still lose some angular momentum. While ETGs with higher stellar mass tend to be slow rotators, PSBs do not follow this trend. We find significant correlations between specific angular momentum and mass-weighted age for the SF and ETG samples, but do not see any significant trends within the short PSB phase. These results indicate that significant evolution in angular momentum must continue to take place as the galaxy ages after the PSB phase. For PSBs to evolve directly into ETGs, they must undergo dry mergers to shed excess angular momentum without causing further epochs of star formation.
post-starburst galaxies — galaxy evolution — observations
software: Astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SDSS-Marvin (Cherinka et al., 2019), Pandas (pandas development team, 2024), Seaborn (Waskom, 2021), SciPy (Virtanen et al., 2020)
Kinematics of PSBsKinematics of Post-starbursts
thanks: Email: bhavya.pardasani@yale.edu
1 Introduction
[…]
In order to trace the evolution of a galaxy, and any mergers undergone by it, we study its kinematics. Resolved integral field unit spectroscopy gives information about the rotational and dispersion components of the stellar velocity, which lets us quantify the specific angular momentum as done by Emsellem et al. (2007), and study the relative significance of stellar rotational velocity ($v$) and stellar velocity dispersion ($\sigma$). The specific angular momentum ($\lambda_{R}$) of a galaxy provides information about whether a galaxy is a slow ($\lambda_{R}\lesssim 0.1$) or a fast ($\lambda_{R}\gtrsim 0.1$) rotator (Emsellem et al., 2007). A comparison of this quantity for PSBs with other galaxies, especially ETGs, can help us connect the merger history with evolution in stellar populations.
The angular momentum content of galaxies is sensitive to their recent merger history because the geometry of the merger and merger mass ratio will affect how much angular momentum is lost or gained. Classifying galaxies kinematically, based on $\lambda_{R}$, can offer valuable insight into their merger history. A loss in angular momentum can happen when two similar mass galaxies merge, in a major merger, along their centers such that they decrease the angular momentum of the resulting galaxy (Naab & Burkert, 2003). A gain in angular momentum could be a product of a merger between a massive galaxy colliding with a significantly lower mass galaxy (a minor merger) with a high value of angular momentum, which leads to the resulting galaxy gaining the excess momentum (Naab & Burkert, 2003). If the merger does not result in a huge influx of cold gas (a dry merger), it can lead to a loss in angular momentum without triggering further star formation.
Observations of galaxies in the immediate aftermath of a strong starburst can be used to test the origins of angular momentum content in ETGs. Major mergers are thought to play a key role in the creation of ETGs (Hopkins et al., 2008), and thus, we would expect the ETGs to be slow-rotating and dispersion-dominated. However, Emsellem et al. (2007, 2011) found that ETGs consist of both slow and fast rotators. Emsellem et al. (2007) suggest that dissipationless (or “dry”) mergers, which cause angular momentum to be expelled outwards, play a key role in the evolution of slow rotators. They propose that multiple channels including dry mergers, gas accretion, and wet mergers likely play a role in the evolution of angular momentum in different classes of ETGs. A key test of how and when slow rotator ETGs lose their angular momentum is the fraction of slow rotators in galaxies in the immediate aftermath of a merger. PSB galaxies appear to be in the middle of evolving from the starburst phase to ETG phase, likely driven by a recent major merger between two gas-rich galaxies (Zabludoff et al., 1996; Yang et al., 2008; Wild et al., 2009; Pawlik et al., 2016; Sazonova et al., 2021), providing a snapshot into the ETG population that may result from this rapid channel of evolution.
Norton et al. (2001), using long-slit spectroscopic data, found 18 out of 20 PSBs from Zabludoff et al. (1996) to be dispersion-dominated, i.e. $v/\sigma<1$, and only 6 out of 20 PSBs exhibited rotation with $v>40$ km/s. Pracy et al. (2009), Swinbank et al. (2012) and Pracy et al. (2013) used integral field spectroscopy to analyze the kinematics of PSBs and found that the PSBs are primarily dominated by fast rotators. Pracy et al. (2012) and Pracy et al. (2013) used the ellipticity-dependent classification presented in Emsellem et al. (2011) to find 4 slow rotators in a sample of 26 E+A galaxies at $z\sim 0.02-0.04$. They found that the slow rotator fraction for the PSBs ($\sim 15-17\%$) is similar to the slow rotator fraction in the ETGs from the ATLAS3D sample ($14\%$) (Emsellem et al., 2011). Similarly, Swinbank et al. (2012) found ${}9\%$ slow rotators in a sample of 11 E+A galaxies at $z=0.06-0.12$. In contrast to findings by Norton et al. (2001) that are more consistent with PSBs experiencing major mergers, results from Swinbank et al. (2012) and Pracy et al. (2013) suggest that their sample of PSBs probably experienced minor mergers. However, a lack of slow rotators in PSBs does not automatically discard the possibility of a major merger as even a 2:1 mass ratio merger can lead to a fast rotator (Emsellem et al., 2011) after triggering a starburst. Moreover, simulated slow rotators usually lose their angular momentum over a period of quite a few Gyrs via dry mergers (Naab et al., 2014). While it is clear that most PSBs are not already slow rotators, their dispersion-dominated kinematics and fraction of slow rotators are similar to ETGs within small number uncertainties. Larger samples of PSBs with spatially-resolved kinematics are needed to test how the angular momentum content of quiescent galaxies evolves over time, and the role of major and minor mergers in the formation of slow rotator ETGs.
In order to further test and compare the results obtained by earlier works, we use integral field spectroscopic data for a large sample of 92 post-starburst galaxies from the Mapping Nearby Galaxies at APO (Apache Point Observatory) (MaNGA Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022) survey, which provides high-resolution spectroscopy for the inner kinematics of 10,000 nearby galaxies. We perform similar $\lambda_{R}$ calculations to Emsellem et al. (2007) for our PSB sample to test whether the PSBs are a result of major mergers, are evolving into ETGs or have undergone various minor/dry mergers. If all of the PSBs have been subjected to equal mass major mergers, we would expect them to have more slow rotators as compared to the ETGs, with perhaps another secular channel producing fast-rotator ETGs. If wet mergers are the main mechanism to produce ETGs, with slow rotators produced in the immediate aftermath of the merger-triggered starburst, we would expect PSBs and ETGs to have a comparable fraction of slow rotators. If minor dry mergers are required to produce slow rotator ETGs, we would expect to see fewer slow rotators among the PSB sample than in the ETG sample.
We present the data and methods in §2. In §3.1, we outline the calculation and corrections performed for the specific angular momentum using the angular momentum expression given by Emsellem et al. (2007). We present our results from our analysis of angular momentum as a function of ellipticity, mass-weighted, light-weighted ages, and asymmetry in §3.2 and §4. Finally, we summarize our results and the implications from this work in §5.
Figure 1: The figure shows the Marvin visualization of the MaNGA data for four of the identified slow rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in Marvin and MaNGA. The middle and right-most panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
Figure 2: The figure shows the Marvin visualization of the MaNGA data for four of the identified fast rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in MaNGA. The middle and right-most panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
2 Data and Target Selection
2.1 Post-starburst Galaxy Selection
There are many methods in the literature for selecting galaxies that have experienced a recent starburst (see French, 2021, for a recent review). This work selects galaxies with a variety of post-starburst ages and AGN properties from a combination of three different samples of PSBs identified in the Sloan Digital Sky Survey (SDSS, York et al. 2000; Strauss et al. 2002), as used in French et al. (2023).
[…]
2.2 MaNGA
The MaNGA Survey (Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022), is one of the three surveys of SDSS-IV that provides data collected for composition and kinematic structure of $\sim$10,000 nearby galaxies via the Sloan 2.5 m telescope at APO. It utilizes optical circular fibers to group 29 fiber bundles of various sizes to create hexagonally-shaped integral field units (IFUs) that are used to spectroscopically observe various sections of a galaxy for a wavelength range of $360-1040$ nm and a spectral resolution of 2000 (Drory et al., 2015; Law et al., 2015; Wake et al., 2017). The fibres have a core diameter of 120 microns which subtends to $2\arcsec$ on the sky. The Sloan telescope has 17 IFUs of various sizes that are used to observe different galaxies simultaneously within a field of view of $3$ deg diametrically ($7$ deg2).
We compared the right ascension and declination of the galaxies from the DR17 version of MaNGA to the $\sim$5000 galaxies from the combined post-starburst sample described in §2.1 to select 92 PSBs [^1] One galaxy is dropped from the sample used by French et al. (2023) because it is not included in the MPA-JHU galSpec analysis, which we use to obtain the stellar mass.. We used NASA-Sloan Atlas (NSA) catalog (Blanton et al., 2011) version v1_0_1 to obtain the photometric properties like Sérsic index (SERSIC_N), ellipticity ($\epsilon=1-$SERSIC_BA), position angle (SERSIC_PHI) and effective radius (SERSIC_TH50) in the r-band, all of which were calculated using a Sérsic fit, for the MaNGA galaxies. We explore the implication of choosing this photometry catalog in Appendix §A.
We obtain the light-weighted (Age_LW_Re_fit) and mass-weighted (Age_MW_Re_fit) ages, and star formation rates (log_SFR_ssp and log_SFR_Ha) for the MaNGA galaxies from SDSS Release 17 pyPipe3D pipeline (v3_1_1), which is a result of the analysis of stellar populations and ionized gas as presented in Sánchez et al. (2016a, b, 2018). Age_LW_Re_fit and Age_MW_Re_fit are the logarithms of the luminosity-weighted and mass-weighted ages of the stellar population in years within the effective radius of the galaxy, respectively. log_SFR_ssp is the logarithm of the integrated star formation rate obtained from the stellar mass that formed in the last 32 Myr. log_SFR_Ha is the logarithm of the integrated star formation rate determined from the integrated H-alpha flux (Sánchez et al., 2016b). We obtain the stellar mass from the “galSpec” galaxy properties from MPA-JHU (Kauffmann et al., 2003; Brinchmann et al., 2004; Tremonti et al., 2004). In order to distinguish the MaNGA SFs from the ETGs, we calculate the specific star-formation rate (sSFR) (SFR/M), where SFR is the integrated star formation rate obtained from the stellar mass log_SFR_ssp. A sSFR $\approx 10^{-11}$ yr-1 approximately divides star-forming and quiescent galaxies (Brinchmann et al., 2004). This sSFR cut is similar to the Halpha EW cut used by French et al. (2018) in selecting PSBs.
Graham et al. (2018) calculated and compared the angular momentum for 2300 galaxies of various types from MaNGA DR14 data, and identified the slow and fast rotators in their sample. We remove the PSBs we identified from their sample, as our angular momentum calculations for the PSBs are based on the MaNGA survey’s DR17 version. The data for the remaining MaNGA star-forming and early-type galaxies is utilized for comparisons and analysis in multiple upcoming plots.
We use data products from the MaNGA data analysis pipeline (Westfall et al., 2019; Belfiore et al., 2019) to measure the angular momentum. We use marvin tools (Cherinka et al., 2019) to obtain the stellar velocity (stellar_vel), stellar velocity dispersion (stellar_sigma), and flux (spx_mflux) for each pixel given as defined by x-coordinate (spx_skycoo: on_sky_x) and y-coordinate (spx_skycoo: on_sky_y). These velocities have been offset to remove their cosmological redshift with respect to the solar barycentric rest frame. To perform this correction, stellar velocities returned by pPXF are used to calculate the observed redshift for individual spaxels, and the effect of the galaxy’s input cosmological redshift is removed to convert the velocities to the reference frame of the galaxy (Westfall et al., 2019). Similarly, the stellar velocity dispersion ($\sigma_{obs}$) is obtained by performing the pPXF fit for the template spectra to the MaNGA data. The $\sigma_{obs}$ needs to be corrected for the effective difference in the instrumental dispersion of the template spectra and MaNGA spectra (Westfall et al., 2019). This correction has been described in Eq 2.
Example data products for the post-starburst galaxies are shown in Figures 1 and 2, which depict the observed galaxy images from SDSS (left-most panels), stellar velocity (middle panels), and stellar velocity dispersion (right-most panels). Figure 1 shows 4 slow rotators and Figure 2 shows 4 fast rotators, from the PSB sample, that were distinguished using the classification discussed in §3.2. These figures showcase the diversity in the post-starburst sample, as galaxies like 11752-3701 (slow rotator) and 9493-1902 (fast rotator), the top galaxies in each figure, have similar morphologies. However, they vary quite a bit in their kinematics, as depicted by the corresponding stellar velocity and dispersion plots. On the other hand, 12071-3702 (slow rotator) and 8080-3704 (fast rotator) are highly elliptical galaxies. The PSB sample also displays a range of merger signatures, like 9507-12704 (fast rotator), and tidal features, like 11753-3704 (fast rotator). These characteristics are quite evident in not just the SDSS images in the left panels, but also from the inner kinematics presented via stellar velocity and stellar velocity dispersion in the middle and right panels.
Name Units Description
Plateifu - Marvin Plate id (Abdurro’uf et al., 2022) to obtain spectroscopic data
MaNGA_Id - MaNGA Id (Abdurro’uf et al., 2022) for the galaxy
RA deg Right ascension (Aihara et al., 2011) of the center of the galaxy
Dec deg Declination (Aihara et al., 2011) of the center of the galaxy
Epsilon - Ellipticity (ϵ=italic-$\epsilon=$1 - SERSIC_BA (Blanton et al., 2011)) of the disk
FWHM_PSF arcsec Full-width half maximum (dispersion)222given by the point spread function which is assumed to be a gaussian (zfwhm)
Sersic_n - Sérsic index (SERSIC_N) (Blanton et al., 2011)
R_eff arcsec Half-light radius (SERSIC_TH50) (Blanton et al., 2011) of the galaxy
Semi_Maj_axis kpc Semi-major axis (= SERSIC_TH50/($1-\sqrt{\epsilon}$) of the galaxy
Phot_angle deg Position angle (SERSIC_PHI) (Blanton et al., 2011) of the disk East of North
Lambda_Re - Corrected angular momentum at half-light radius
Stellar_mass M Stellar mass from galSpec MPA-JHU (Kauffmann et al., 2003)
SFR_ssp M/yr star formation rate (log_SFR_ssp) (Sánchez et al., 2016a, b, 2018) of the galaxy
SFR_ha M/yr star formation rate (log_SFR_Ha) (Sánchez et al., 2016a, b, 2018) of the galaxy in H-alpha
LW_age Gyr Light-weighted age (Age_LW_Re_fit) (Sánchez et al., 2016a, b, 2018) of the galaxy
MW_age Gyr Mass-weighted age (Age_MW_Re_fit) (Sánchez et al., 2016a, b, 2018) of the galaxy
Asymmetry - Morphological asymmetry (A) (Hernández-Toledo et al. (2010); Cheng et al. (2011); Vázquez-Mata et al. (2022))
of the galaxy
Table 1: PSB galaxy properties. Full table available online.
2.3 Other sources of data
We obtain the data for a different sample of 260 early-type galaxies from the ATLAS3D Project (Cappellari et al., 2011). This project obtained its data from the SAURON integral-field unit on the William Herschel Telescope (WHT), combined with deep imaging done with MegaCam on the Canada-France-Hawaii Telescope (CFHT).
We use data sets provided by multiple papers under the umbrella of the ATLAS3D project to obtain parameters that enable us to make comparisons for angular momentum in the context of luminosity-weighted (LW) age, mass-weighted (MW) age, and stellar mass. We acquire ellipticity (Epsilon_e) and angular momentum (LambdaR_e) at effective radius from Emsellem et al. (2011). Emsellem et al. (2011) calculated the $\lambda_{Re}$ using the expression from Emsellem et al. (2007), which is given in Eq 1.
[…]
Name Units Description
Galaxy - Name of the galaxy (Emsellem et al., 2011)
Lambda_obs - Specific angular momentum as calculated by Emsellem et al. (2011)
Epsilon - Ellipticity (Epsilon_e) from (Emsellem et al., 2011)
FWHM_PSF arcsec Full-width half maximum (dispersion)333used an average value of 1.25” (Scott et al., 2013)
Sersic_n - Sérsic index (n) (Krajnović et al., 2013) of the galactic disk
Semi_Maj_axis kpc Semi-major axis calculated using logRem̂aj and D (Cappellari et al., 2013a) of the galaxy
Lambda_corr - Lambda_obs corrected for beam-smearing (Graham et al., 2018)
Stellar_mass M Stellar mass calculated using logLum (Cappellari et al., 2013a)
and logML_star (Cappellari et al., 2013b) as described in §2.3
LW_age Gyr Light-weighted age (Age_SSP) (McDermid et al., 2015) of the galaxy
MW_age Gyr Mass-weighted age (Age_SHF) (McDermid et al., 2015) of the galaxy
Table 2: ETG properties for the ATLAS3D comparison sample. Full table available online.
3 Results
3.1 Angular momentum calculation and correction
Figure 3: Angular momentum at half-light radius, $\lambda_{Re}$ vs. ellipticity at half-light radius, $\epsilon_{Re}$, for 92 PSBs from the MaNGA Survey and 260 ETGs from the ATLAS-3D Project, with point sizes scaled by log stellar mass. The blue and red contour lines represent the 1577 SFs and 964 ETGs from MaNGA respectively. The black dashed line at $\lambda_{Re}=0.1$ separates the slow and fast rotators according to the classical definition, while the black solid line given by $0.31\times\sqrt{\epsilon_{Re}}$ (Emsellem et al., 2011) performs a similar function as the black dashed line, but takes into account the ellipticity of the galaxies in the classification of slow- and fast rotators. We find that a higher fraction of ETGs are slow rotators than the PSB and SF samples, with the SF samples having significantly higher average $\lambda_{Re}$.
We calculate the specific angular momentum for each galaxy using the expression from Emsellem et al. (2007) as:
$\lambda_{R}=\frac{\sum_{i=1}^{N_{p}}F_{i}R_{i}|V_{i}|}{\sum_{i=1}^{N_{p}}F_{i}R_{i}\sqrt{{V_{i}}^{2}+{\sigma_{i}}^{2}}}$ (1)
where $F_{i}$, $R_{i}$, $V_{i}$, and $\sigma_{i}$ are the flux, distance from the center, stellar velocity, and dispersion respectively in the $i$th pixel. For the MaNGA data, we utilize the Python functions provided by Graham et al. (2018).
We calculate the angular momentum (given in column 3 of Table 1) of the 92 PSBs using Eq 1 and the lambdaR_e_calc function defined by Graham et al. (2018) by passing the stellar velocity, stellar velocity dispersion and flux in each pixel (§2). We also provide effective radius, ellipticity, position angle, FWHM_psf, and Sérsic index (all of which are described in Table 1) from the NSA catalog, in order to determine the 1 $R_{e}$ ellipse in which to calculate $\lambda$. The stellar velocity dispersion has been corrected for the instrumental dispersion (stellar_sigmacorr) using the given correction provided by Westfall et al. (2019):
$\sigma_{corrected}=\sqrt{\sigma_{raw}^{2}-\sigma_{corr}^{2}}$ (2)
According to Graham et al. (2018), the finite point-spread function (PSF) results in the smearing of the velocity field. To correct for this effect, they provide an optional analytic correction as part of their lambdaR_e_calc function. This correction can be applied to regular rotators whose semi-major axis is larger than the PSF dispersion (sig_PSF) and Sérsic index is in the range $0.5-6.5$. In order to apply this correction to all the galaxies in our sample, we replace the Sérsic indices with 0.5001 if $n<0.5$ and with 6.4999 if $n>6.5$. Thus, we obtain the beam-corrected angular momentum for the 92 PSBs. All the parameters used and the angular momentum calculated for the PSB are presented in Table 1.
[…]
3.2 Identifying slow and fast rotators
We plot the beam-corrected specific angular momentum at half-light radius for 92 MaNGA PSBs and 260 ETGs from ATLAS3D against the ellipticity at half-light radius in Figure 3. The points are scaled in size and color by stellar mass (acquired as described in §2.1 §2.3 for PSBs and ETGs respectively), with the corresponding mass given in the legend. The blue and red contours depict the 1577 star-forming and 964 early-type MaNGA galaxies respectively whose data is obtained from Graham et al. (2018). We find that a higher fraction of ETGs are slow rotators than the PSB and SF samples, with the SF samples having significantly higher average $\lambda_{Re}$.
As shown in Figure 3, we show two criteria to separate the slow and fast rotators: the classical $\lambda_{Re}=0.1$ (black dashed line) boundary condition and the curve given by $0.31\times\sqrt{\epsilon_{Re}}$ (black dotted line) (Emsellem et al., 2011). Emsellem et al. (2011) found that the specific angular momentum generally increases with the ellipticity, even for non-regular rotators, so they developed a relationship between $\lambda_{R}-\epsilon$ to distinguish slow and fast rotators. This relationship also accounts for the increasing galactic anisotropy with ellipticity. Thus, $\lambda_{Re}=0.31\times\sqrt{\epsilon_{Re}}$ is a better measure of whether a galaxy is a slow or fast rotator. Using this as the preferred limiting condition, we find that 5 out of 92 ($\approx$5%) PSBs, 36 out of 260 ($\approx$14%) ETGs from ATLAS3D, 64 out of 1577 ($\approx$4%) MaNGA SFs, and 213 out of 964 ($\approx$22%) MaNGA ETGs are slow rotators. We also find that $84\%$ of SFs have higher angular momentum than the maximum $\lambda_{Re}$ of the PSB sample. Thus, the ETGs generally have lower angular momentum as compared to the PSBs, which tend to rotate slower than the SFs.
The scaling and coloring of points based on stellar mass shows that galaxies, especially ETGs, with higher stellar mass prefer to rotate slower as compared to their counterparts. We explore trends in angular momentum with stellar mass in Figure 4. In this Figure, we can see a distinct population of high stellar mass slow rotators in the ETG sample, that is not present for the PSBs and SF galaxies. To quantify this observation, we consider the fraction of slow rotators in the top 30th percentile of stellar mass for each galaxy type. $42\%$ of MaNGA SFs, $40\%$ of MaNGA PSBs, $58\%$ of MaNGA ETGs and $58\%$ of ATLAS3D ETGs that are slow rotators, are in the top 30th percentile of stellar mass. While the slow rotator SF and PSB galaxies are slightly over-represented in the highest stellar mass galaxies, the slow rotator ETGs are significantly more biased towards the higher mass galaxies. We find similar levels of over-representation when we consider the top 10th or 20th percentiles of stellar mass.
We consider the fraction of slow rotators in the PSBs within the different parent samples. Of the 92 PSBs, there are 25 E+As, 19 SPOGs, and 77 PCAs (with significant overlap). Two of the 5 slow rotators are in multiple classifications, and we find 1 E+A, 1 SPOG, and 5 PCA slow rotators. Within the small number uncertainties, the fraction of slow rotators is consistent between each parent PSB sample.
Figure 4: Angular momentum vs. stellar mass for 92 PSBs (black scatter), 1577 SFs (blue contours) and 964 ETGs (red contours) from MaNGA, and 260 ATLAS3D ETGs (red scatter). We find that galaxies, especially the early-types, with higher stellar mass tend to rotate slowly as compared to galaxies with lower stellar mass. While most of the slow rotator ETGs are amongst the most massive ETGs, a similar effect is not seen for the PSB and SF samples. We explore this effect quantitatively in the text. This implies that the high stellar mass slow rotator ETGs have experienced further growth and loss of angular momentum since they may have passed through the PSB phase.
4 Discussion
4.1 Evolution of angular momentum with age
Figure 5: Left: Angular momentum at half-light radius, $\lambda_{Re}$, vs. mass-weighted age for 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for PSBs. We observe a clear distinction in age between the SF and ETG samples, with PSB galaxies having the youngest MW ages. A significant anti-correlation between MW age and $\lambda_{Re}$ is seen for the SF and ETG samples, though not within the PSB sample. At mass-weighted ages $\sim$5 Gyr, where the oldest PSBs have similar mass-weighted ages to the youngest ETGs, the PSB and ETG distributions in angular momentum overlap. For older ETGs, we see that a trend towards lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. Right: Angular momentum at half-light radius, $\lambda_{Re}$, vs. light-weighted age 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. Similar to Figure 5a, we observe a clear distinction between the SF and ETG samples, though now in the case of LW age, the PSB ages are typically between the SF and ETG samples. For the LW ages, we observe a significant trend within the ETG samples, but a weaker trend in the SF sample and no significant trend in the PSB sample. The fact that the angular momentum is more closely coupled with the mass-weighted age implies that it is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age.
To further our understanding of the evolution of angular momentum across the three classes, we study its evolution with the age of the stellar populations. To do so, we plot the angular momentum against mass- and luminosity-weighted age (as shown in Figure 5. As discussed in §2.2, we obtain the respective ages for MaNGA SFs, PSBs and ETGs from SDSS Release 17 pyPipe3D pipeline (v3_1_1), which is a result of the analysis of stellar populations and ionized gas as presented in Sánchez et al. (2016a, b, 2018).
We observe that angular momentum decreases with increasing age across all three samples for MW and LW age, with the early types having older MW and LW ages and more slow rotators than PSBs and SFs. At mass-weighted ages $\sim$5 Gyr, where the oldest PSBs have similar mass-weighted ages to the youngest ETGs, the PSB and ETG distributions in angular momentum overlap. For older ETGs, we see that a trend towards lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. To quantify this, we compute the Spearman correlation (presented in Table 3) for all three classes of galaxies for MW and LW age. We find significant correlations of angular momentum with mass-weighted age for the SF and ETG samples, though not within the PSB sample. For the light-weighted ages, we see a weaker correlation for the SF sample. The fact that the angular momentum is more closely coupled with the mass-weighted age implies that it is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age.
We also test whether we see evolution within the PSB phase by testing the correlation between angular momentum and age since the starburst ended using the stellar population fits by French et al. (2018). Consistent with the lack of correlation in MW or LW age, we observe no significant trend within the PSB in angular momentum vs. starburst age.
Table 3: Spearman result for MW and LW ages for MaNGA SFs, PSBs, and ETGs
Galaxy MW age LW age
SFs ($\rho$ = -0.22, p $\approx 10^{-16}$) ($\rho$ = -0.06, p = 0.04)
PSBs ($\rho$ = -0.11, p = 0.31) ($\rho$ = -0.16, p = 0.13)
ETGs ($\rho$ = -0.33, p $\approx 10^{-24}$) ($\rho$ = -0.29, p $\approx 10^{-19}$)
4.2 Evolution of angular momentum with asymmetry
Figure 6: Angular momentum at half-light radius, $\lambda_{Re}$, versus asymmetry, $A$, for 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. The ETGs tend to be fast and slow rotators for similar asymmetry values, while angular momentum is correlated with asymmetry for the SFs. The PSBs lie between the ETGs and SFs in both angular momentum and asymmetry.We do not observe a significant trend between angular momentum and asymmetry within the PSB sample. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
Asymmetric features in galaxies can be used to trace their recent merger history due to the presence of tidal tails, shells, and companion galaxy remnants, as well as to separate galaxies between morphological types. The rotational asymmetry $A$, a purely morphological parameter, is calculated by rotating a galaxy by 180o and subtracting it from the original image, and further applying relevant corrections to take into account background noise etc.
PSBs tend to exhibit high asymmetry compared to the ETGs, which could be due to tidal/relic features associated with the recent starburst and/or recent interaction with another galaxy or dust lanes in the case of dusty PSBs (Yang et al., 2008; Sazonova et al., 2021). The SFs have high $A$ due to the presence of spiral arms and flocculence in the disk. However, the ETGs tend to have low $A$ due to the absence of tidal features or spiral arms.
The asymmetry, thus, provides a way to test for the effects of recent mergers and interaction in galaxies and test for trends with angular momentum content. We would expect the PSBs resulting from major mergers to be characterized by significantly lower angular momentum, however, the expected trends depend on the relative timescales for tidal features to be present and angular momentum to be lost. Low values of $A$ could be present in post-major-merger galaxies due to the fact that their merger features fade quickly as the PSBs age (Pawlik et al., 2016). The PSBs also have a large range of angular momentum, which can obscure any subtle trends we might observe.
Asymmetry values for the PSBs, ETGs, and SFs in the MaNGA survey are obtained from the MaNGA Visual Morphology Catalogue (Vázquez-Mata et al., 2022). Following digital image post-processing, the catalogue detects the inner structures and external, low surface brightness characteristics. The morphological classification process was implemented empirically, based on the methods described in Hernández-Toledo et al. (2010) and Cheng et al. (2011). They reported an estimation of the CAS parameters (Concentration, Asymmetry, and Clumpiness) derived from the Legacy Surveys and SDSS imaging, following the method by Conselice (2003).
Figure 6 compares the angular momentum against rotational asymmetry for 92 PSBs, 1357 SFs, and 903 ETGs from the MaNGA survey. The black line is the running average for the PSBs. We observe that the slow and fast rotators in the ETG sample are concentrated around small values of asymmetry, however, the fast rotators in the SF sample tend to exhibit high asymmetry values. The finding of higher angular momentum in the SF galaxies with more asymmetry may be due to a similar effect as seen by Romanowsky & Fall (2012). They find that later Hubble-type galaxies and galaxies with lower bulge-to-total light ratios have more angular momentum, even within the class of spirals. The PSBs, on the other hand, do not seem to exhibit significant correlation (Spearman correlation: $\rho$ = 0.19, p = 0.11) for angular momentum and asymmetry. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
Deeper photometry may reveal more subtle trends between the asymmetry caused by merger features and the effect on angular momentum. Rutherford et al. (2024) used deep imaging from Subaru-Hyper Suprime Cam to test for the presence of tidal features and shells around galaxies as a function of angular momentum, finding a correlation between the presence of shells and lower angular momentum in a subset of the younger galaxies. Deeper photometry in next-generation surveys such as the Rubin Observatory LSST will provide useful constraints on the correlation of merger features with angular momentum during the PSB phase and into the ETG phase.
5 Conclusions
We trace the evolution of a galaxy as it evolves from the post-starburst phase to the quiescent phase by comparing and contrasting spatially resolved properties of 92 post-starbursts from the MaNGA survey with 260 early-type galaxies from the ATLAS3D project, 1577 MaNGA star-forming galaxies, and 964 early-types from MaNGA. We attempt to glean information about the merger history of these galaxies by studying their inner kinematics and classifying them as slow and fast rotators based on their angular momentum and ellipticity at the effective radius. Our conclusions are as follows:
  • (i)
    We find that the PSB sample has fewer slow rotators than the ETG samples from both ATLAS3D and the MaNGA DR14 (Graham et al., 2018), with the PSB sample consisting of $\approx 5\%$ (5/92), the ATLAS3D ETG sample having $\approx 14\%$ (36/260), SFs from MaNGA having $\approx 4\%$ (64/1577), and MaNGA ETGs having $\approx 22\%$ (213/964) slow rotators. Star-forming galaxies from Graham et al. (2018) have consistently high values of angular momentum, with a mean value above even the fastest rotator PSBs in our sample.
  • (ii)
    We note that the angular momentum is generally lower for ETGs with higher stellar mass, which is not the case for the PSBs or SFs. This is evident from our finding that $\sim 58\%$ of slow rotator MaNGA ETGs and $58\%$ of slow rotator ATLAS3D ETGs are in the top 30th percentile of the respective ETG stellar mass distributions, while only $40\%$ slow-rotating MaNGA PSBs and $42\%$ slow-rotating MaNGA SFs are in the top 30th percentile of the respective PSB and SF stellar mass distribution.
  • (iii)
    From our analysis of the angular momentum in the context of mass-weighted and light-weighted ages, we find that the angular momentum of SFs and ETGs is more tightly correlated with MW age as compared to the PSBs. We observe a significant decrease in angular momentum with MW age for both the ETG and SF samples, though not within the PSB sample. For LW age, we only observe a significant correlation within the ETG sample.
  • (iv)
    Our analysis of galaxy kinematics with respect to asymmetry does not suggest a clear trend for the PSBs (Spearman correlation: $\rho$ = 0.19, p = 0.11). Both slow and fast rotator ETGs have similarly low values of asymmetry, while the SF sample displays a correlation of higher asymmetry in the highest angular momentum galaxies.
This work constrains the possible formation and evolution of ETGs. Roughly half of ETGs are expected to have gone through a PSB phase after rapidly quenching (Wild et al., 2020), so if slow rotator ETGs are created through gas-rich merger mergers, we should observe a similar or greater fracion of slow rotators in the PSBs. However, we find more slow rotators in the ETG sample as compared to the PSBs. The tight correlation of angular momentum with the mass-weighted age, as compared to the luminosity-weighted age, indicates that angular momentum is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age. PSBs already have lower angular momentum than SF/disk galaxies, and since kinematic evolution takes place over the many-Gyr lifetime of a galaxy, PSBs must still lose significant angular momentum before evolving to ETGs. This is consistent with the trends in angular momentum with stellar mass, which imply that the slow rotator ETGs have gained mass and lost angular momentum over many Gyr since they last formed stars in situ or experienced a starburst. The long timescale of kinematic evolution is probably why we do not observe significant trends for angular momentum with asymmetry as the asymmetry varies quite rapidly during/after a merger. These implications suggest that the PSBs will undergo multiple major/minor dry merger events as they evolve to the ETG phase while shedding off their angular momentum. This scenario is consistent with the findings of simulations presented in Naab et al. (2014), where the slow rotator ETGs tended to have many dry mergers over being quiescent for a long time.
Acknowledgements
We thank Akshat Tripathi for useful discussions. This work was made possible through the Preparing for Astrophysics with LSST Program, supported by the Heising-Simons Foundation and managed by Las Cumbres Observatory.
Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org.
SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix A Comparison of Photometry
In order to analyze the kinematics within 1 Re, we need to know the shape and size of the galaxy. We also need the Sérsic index to estimate the impact of beam smearing on the photometry. We compare the photometric properties from the NSA catalog with the ones given in Table 3 of Simard et al. (2011). Simard et al. (2011) performed bulge+disk decompositions in g- and r-band by convolving the point spread function in 2D for 1,123,718 galaxies from the Legacy portion of SDSS Release 7. They use three fitting models - a pure Sérsic model, an nb = 4 bulge + disk model, and a Sérsic (free nb ) bulge + disk model. We chose the data calculated using the pure Sérsic model as NSA also used a Sérsic fit, making the two sets of photometric properties comparable. We compare the galaxy Sérsic index (ng), the galaxy position angle (phi), galaxy ellipticity (e), and the galaxy half-light radius in r-band (Rhlr).
To test the consistency of the photometry across all the galaxies, and not just the PSBs, we plot the photometric properties from the NSA catalog against the respective properties from Simard et al. (2011), as shown in Figure 7. The gray contours represent all the galaxies in both catalogs, and the blue contour lines represent the 92 PSBs across the two catalogs. We observe that the scatter in the photometric properties of the PSBs around the red 1:1 line is quite consistent with the overall scatter of all the galaxies for all the properties. This means that overall, we do not expect a particular bias for the PSBs, although the scatter in properties can be large.
To further test our choice to use the NSA photometry, we compare the NSA catalog to the MaNGA PyMorph DR17 photometric catalogue. PyMorph provides photometric properties determined by fitting Sérsic and Sérsic+Exponential models to the 2D surface brightness profiles of galaxies in the MaNGA DR17 release (Meert et al., 2013; Fischer et al., 2017; Bernardi et al., 2017; Domínguez Sánchez et al., 2021). PyMorph was re-run to incorporate the SDSS DR14 images, improved bulge-to-disk decomposition, and all the fits were re-fit if deemed necessary (Fischer et al., 2018; Domínguez Sánchez et al., 2021). Since NSA performed a pure Sérsic fit, we selected the PyMorph photometry obtained from fitting the Sérsic model. We compare the Sérsic index (N_S), photometric angle (90 - PA_S - SPA), ellipticity (1-BA_S), and effective radius (obtained from the semi-major axis, A_hl_S) in r-band from PyMorph runs with the corresponding properties in r-band from the NSA catalog, as shown in Figure 8. The gray contours represent all the galaxies in both catalogs, and the blue contour lines represent the 92 PSBs across the two catalogs. Unlike the case for the Simard et al. (2011) parameters, the scatter in the PSBs is larger than the typical scatter for normal galaxies, which indicates the parameters derived for the PSBs may contain significant outliers.
Figure 7: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the properties given in Table 3 of Simard et al. (2011), all of which were determined using a pure Sérsic model. The gray shaded contour region corresponds to all the galaxies while the blue contour lines represent the 92 PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. We observe that the scatter in the photometric properties of the PSBs around the red 1:1 line is quite consistent with the overall scatter of all the galaxies for all the properties. This means that overall, we do not expect a particular bias for the PSBs, although the scatter in properties can be large.
The scatter observed in photometry properties between these catalogs may mean that the choice of catalog influences our measurement of specific angular momentum. The $\lambda_{Re}$ obtained using the three catalogs are plotted against each other in Figure 9. We do not observe any particular trend in the scatter and instead find that the points tend to scatter around the 1:1 black line, proving that choosing another catalog would not affect our results and conclusions drastically. We find 5.4% (5/92) slow rotators using the NSA catalog, 7.6% (7/92) slow rotators using Simard et al. (2011) photometric properties, and 9.8% (9/92) slow rotators using the PyMorph catalogue. In all three cases, the PSBs have significantly fewer slow rotators than the ETG samples.
As a further test, we over-plot the SDSS images of the PSBs with ellipses created using the photometric properties of the respective catalogs. We observe that the ellipse corresponding to NSA photometry aligns quite well with the galaxy, however, the ellipses traced by photometric properties from Simard et al. (2011) and the PyMorph catalogue do not align as well, and in some cases are aligned away from the galaxy by $\sim 90^{o}$ for Simard et al. (2011) photometry. Thus, we use the NSA catalog to measure $\lambda_{Re}$ in this work as described in §3.1.
Figure 8: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the corresponding pure Sérsic model properties provided in MaNGA PyMorph DR17 photometric catalogue. The gray shaded contour region corresponds to all the galaxies while the blue contour lines represent the 92 PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. Unlike the case for the Simard et al. (2011) parameters, the scatter in the PSBs is larger than the typical scatter for normal galaxies, which indicates the parameters derived for the PSBs may contain significant outliers.
Figure 9: Comparison of the angular momentum obtained using photometry from the NSA catalog, Simard et al. (2011) and MaNGA PyMorph DR17 photometric catalogue for the PSBs. We do not observe a systematic scatter as the points tend to scatter about the 1:1 black line. We find 5.4% (5/92) slow rotators using the NSA catalog, 7.6% (7/92) slow rotators using Simard et al. (2011) photometric properties, and 9.8% (9/92) slow rotators using the PyMorph catalogue. In all three cases, the PSBs have significantly fewer slow rotators than the ETG samples.